OpenStudy (anonymous):

Find the foci of the hyperbola defined by this equation: ((x-4)^2)-(1)-((y+4)^2)-(9)=1

OpenStudy (mertsj):

\[\frac{(x-4)^2}{1}-\frac{(y+4)^2}{9}=1\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

\[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\]

OpenStudy (mertsj):

That is the equation of a hyperbola with center (h,k)

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

a^2+b^2=c^2

OpenStudy (mertsj):

So identify a^2 and b^2 and use them to find c because the foci are c units to the right and c units to the left of the center.

OpenStudy (anonymous):

I don't know how to do that....

OpenStudy (mertsj):

What is a^2?

OpenStudy (anonymous):

1^2

OpenStudy (mertsj):

The number in your equation that replaces a^2 is 1

OpenStudy (mertsj):

What is b^2?

OpenStudy (anonymous):

9

OpenStudy (mertsj):

Now take the equation I posted. Replace a^2 with 1. Replace b^2 with 9. Solve for c

OpenStudy (anonymous):

10?

OpenStudy (mertsj):

c^2=10 so c is sqrt10

OpenStudy (mertsj):